trouble here!! help please
find the indicated outputs for f(x) = 5x^2-4x. Can you explain this to me step by step, much appreciated

Answer by RAY100(1637)

y= 5x^2 -4x = x(5x-4)=0,,,,factor
,
x=0,,,,x=4/5,,,,
,
(0,0),,,,,(4/5,0),,,,are roots,,,,zeros
.